Exercise 3-3 Compute ABC Product Costs [LO3-3] Larner Corporation is a diversified manufacturer of industrial goods. The company's activity-based costing system contains the following six activity cost pools and activity rates: Activity Cost Pool Activity Rates Labor-related $ 7.00 per direct labor-hour Machine-related $ 3.00 per machine-hour Machine setups $ 40.00 per setup Production orders $ 160.00 per order Shipments $ 120.00 per shipment General factory $ 4.00 per direct labor-hour Cost and activity data have been supplied for the following products: J78 B52 Direct materials cost per unit $ 6.50 $ 31.00 Direct labor cost per unit $ 3.75 $ 6.00 Number of units produced per year 4,000 100 Total Expected Activity J78 B52 Direct labor-hours 1,000 40 Machine-hours 3,200 30 Machine setups 5 1 Production orders 5 1 Shipments 10 1 Required: Compute the unit product cost of each product listed above. Unit Product cost for: J78 and B52
Answer #1

Calculate Unit product Cost

J78 B52
Labor related 1000*7 = 7000 40*7 = 280
Machine related 3200*3 = 6300 30*3 = 90
Machine setup 40*5 = 200 40*1 = 40
Production order 160*5 = 800 160*1 = 160
Shipments 120*10 = 1200 120
General factory 1000*4 = 4000 40*4 = 160
Total overhead 19500 850
Unit 4000 100
Overhead cost per unit 4.875 8.5
Direct material 6.50 31
Direct labor 3.75 6.0
Unit product cost 15.125 45.50